OpenStudy (anonymous):

Use the functions to answer the question. f(x) = 4x -1 ; g(x) = 5x^2 Which is equal to (f times G) (x) (Points : 4) Option A: 80x^2 - 40x +5 Option B: 20x^3 -5x^2 Option C: 5x^2 + 4x -1 Option D: 20x^2 -1

OpenStudy (anonymous):

f(g(x)) would look like this. 4(5x^2)-1. You solve for x.

OpenStudy (anonymous):

How do I solve for x?

OpenStudy (anonymous):

20x^2-1. You don't solve for x. My mistake. the answer would be 20x^2-1

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

f(x) =x^2 -4 ; g(x) = x+ 2 which is equal to (f + g) (x) x^2 -2 x^2 + 4x x^2 + x -2 x^3 + 2x^2 -4x -8